OpenStudy (help_people):

algebra 1 help please?

OpenStudy (help_people):

The following graph describes function 1, and the equation below it describes function 2: Function 1 graph of function f of x equals negative x squared plus 8 multiplied by x minus 15 Function 2 f(x) = −x2 + 2x − 3 Function ____ has the larger maximum. (Put 1 or 2 in the blank space) Numerical Answers Expected!

OpenStudy (xapproachesinfinity):

parabola has the maximum at the vertex if a<0 right

OpenStudy (xapproachesinfinity):

agree

OpenStudy (xapproachesinfinity):

yes , no? what do you think

OpenStudy (xapproachesinfinity):

i hate when the poster just sits there waiting for an answer

OpenStudy (help_people):

srry internet connecton

OpenStudy (help_people):

and i agree (/yes)

OpenStudy (help_people):

@xapproachesinfinity

OpenStudy (xapproachesinfinity):

good

OpenStudy (xapproachesinfinity):

so what are the coordinates of any parabola's vertex

OpenStudy (xapproachesinfinity):

do you know?

OpenStudy (xapproachesinfinity):

\(\Large (\frac{-b}{2a},f(\frac{-b}{2a}))\) this is the vertex of any parabola

OpenStudy (xapproachesinfinity):

if a<0 f(-b/2a) is the maximum if a>0 f(-b/2a) is minimum

OpenStudy (xapproachesinfinity):

@help_people

OpenStudy (help_people):

no i dont know @xapproachesinfinity